OpenStudy (anonymous):

3x-6y=6 ....graph the linear function by finding x and y intercepts. do i just substitute x and y for 0? i don't understand. thanks for any help!!

OpenStudy (anonymous):

Well, I think the easiest way would be to put the equation in slope intercept form. (y = mx + b), can you get y by itself?

OpenStudy (anonymous):

Yes. -6y=-3x+6

OpenStudy (anonymous):

Ok, now get y completely by itself. Divide everything by -6.

OpenStudy (anonymous):

would it be y=3/6x-1?

OpenStudy (anonymous):

Perfect. 3/6 = 1/2. Reduce your fractions ;)

OpenStudy (anonymous):

So in slope intercept form the "b" part is the y-intercept. Because when x is 0, you can see the (1/2x) just goes away completely. So what's the y-intercept?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

You got it.

OpenStudy (anonymous):

So to find the x-intercept. Set y = 0. Then solve for x. 0 = (1/2)x - 1

OpenStudy (anonymous):

you set y for 0 in the original problem?

OpenStudy (anonymous):

No, set 0 in for y in the "slope intercept" form.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the slope is 1/2 and the y is -1?

OpenStudy (anonymous):

one more question. how do you write the equation by using function notation?

OpenStudy (anonymous):

In this sense, y is the same as f(x) So simply: f(x) = (1/2)x - 1 I want to be sure you found the x-intercept.

OpenStudy (anonymous):

okay i got it. thank you! that was the write answer.

OpenStudy (anonymous):

Ok, good deal.

OpenStudy (anonymous):

thanks again for all the help!!!!! See ya!

OpenStudy (anonymous):

No problem.